# Highly edge-connected subgraphs

### v1

@@ Abstract
    $G$ has minimum degree at least (3s1)+(3t1)+1(3s-1)+(3t-1)+1. By a result of Stiebitz [7], $V(G)$ has a partition into two sets $A^{\prime}$ and $B^{\prime}$ such that $G[A^{\prime}]$ has minimum degree at least 3s13s-1 and $G[B^{\prime}]$ has minimum degree at least 3t13t-1. By Theorem On highly (vertex-)connected subgraphs, $G[A^{\prime}]$ has an $(s+1)$-connected subgraph on some vertex set $A$ of size >2s>2s and $G[B^{\prime}]$ has a $(t+1)$- connected subgraph on some vertex set $B$ of size >2t>2t. We now ignore $A^{\prime}$ and $B^{\prime}$.
    Among all disjoint $A,B\subseteq V(G)$ with $|A|>2s$ and $|B|>2t$ such that $G[A]$ is $(s+1)$-connected and $G[B]$ is $(t+1)$-connected, we choose $A$ and $B$ with minimum size of $X=V(G)\setminus(A\cup B)$. Assume $X\neq\emptyset$.
    Then $G[A\cup X]$ is not $(s+1)$-connected, that is, there is some $S\subseteq A\cup X$ of at most $s$ vertices such that $G[(A\cup X)\setminus S]$ decomposes into more than one connected component. One of them has to be disjoint from $A$. Let $Y\subseteq X$ be its vertex set.
    Now $G[B\cup Y]$ is not $(t+1)$-connected, that is, there is some $T\subseteq B\cup Y$ of at most $t$ vertices such that $G[(B\cup Y)\setminus T]$ decomposes into more than one connected component. One of them has to be disjoint from $B$. Let $Z\subseteq Y$ be its vertex set. The neighborhood of $Z$ is contained in $S\cup T$ of size at most $s+t$, a contradiction to the $(s+t+1)$-connectedness of $G$.
@@
    By a result of Kaneko [8] on triangle-free graphs, $V(G)$ has a partition into two sets $A^{\prime}$ and $B^{\prime}$ such that $G[A^{\prime}]$ has minimum degree at least $2s$ and $G[B^{\prime}]$ has minimum degree at least $2t$. Now use Theorem On highly (vertex-)connected subgraphs to obtain disjoint $(s+1)$- and $(t+1)$-connected subgraphs. Then proceed as before.
@@
    By Theorem On highly edge-connected subgraphs and Theorem On highly edge-connected subgraphs for small $k$, there exists a function $g(k)=\big{(}1+o(1)\big{)}k$,$g(k)<2k$ such that for every $k$, every graph of minimum degree at least $g(k)$ contains a $(k+1)$-edge-connected subgraph. We choose $f(r)=\max\big{\{}g(r-s)+g(s)+1\,|\,1\leq s\leq\frac{r}{2}\big{\}}$. We have that $f(r)=\big{(}1+o(1)\big{)}r$ for $r\to\infty$:
    If $s\geq\sqrt{r}$, then $g(r-s)+g(s)+1=\big{(}1+o(1)\big{)}(r-s)+\big{(}1+o(1)\big{)}s+1=\big{(}1+o(1)\big{)}r$.Otherwise, $g(r-s)+g(s)+1<\big{(}1+o(1)\big{)}(r-s)+2s+1\leq\big{(}1+o(1)\big{)}r+o(r)=\big{(}1+o(1)\big{)}r$.
    Now let $G$ be an $(s+t+1)$-edge-connected graph of minimum degree at least $f(s+t)\geq g(s)+g(t)+1$. Again by Stiebitz [7], $V(G)$ has a partition into two sets $A^{\prime}$ and $B^{\prime}$ such that $G[A^{\prime}]$ and $G[B^{\prime}]$ have minimum degree at least $g(s)$ and $g(t)$, respectively. We obtain disjoint $(s+1)$- and $(t+1)$-edge-connected subgraphs. Since Thomassen’s argument also works for edge-connectedness, we may proceed as before. $\Box$

### v2

@@ Abstract
    $G$ has minimum degree at least $(3k_{1}-1)+(3k_{2}-1)+1$. By a result of Stiebitz [8], $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has minimum degree at least $3k_{i}-1$. By Theorem On highly connected subgraphs, $G[A_{i}]$ has a $(k_{i}+1)$-connected subgraph on some vertex set $B_{i}$ of size $2k_{i}$, for $i\in\{1,2\}$. We choose maximal disjoint $C_{1},C_{2}\subseteq V(G)$ such that $C_{i}\supseteq B_{i}$ and $G[C_{i}]$ is $(k_{i}+1)$-connected, for $i\in\{1,2\}$.
    Assume that $X=V(G)\setminus(C_{1}\cup C_{2})\neq\emptyset$. Then $G[C_{1}\cup X]$ is not $(k_{1}+1)$-connected, that is, there is some $S_{1}\subseteq C_{1}\cup X$ of at most $k_{1}$ vertices such that $G[(C_{1}\cup X)\setminus S_{1}]$ has more than one connected component. One of them has to be disjoint from $C_{1}$. Let $Y\subseteq X$ be its vertex set.
    Now $G[C_{2}\cup Y]$ is not $(k_{2}+1)$-connected, that is, there is some $S_{2}\subseteq C_{2}\cup Y$ of at most $k_{2}$ vertices such that $G[(C_{2}\cup Y)\setminus S_{2}]$ has more than one connected component. One of them has to be disjoint from $C_{2}$. Let $Z\subseteq Y$ be its vertex set. The neighborhood of $Z$ is contained in $S_{1}\cup S_{2}$ of size at most $k_{1}+k_{2}$, a contradiction to the $(k_{1}+k_{2}+1)$-connectedness of $G$.
@@
    By a result of Kaneko [9] on triangle-free graphs, $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has minimum degree at least $2k_{i}$. Now use Theorem On highly connected subgraphs to obtain disjoint $(k_{1}+1)$- and $(k_{2}+1)$-connected subgraphs. Then proceed as before.
@@
    By a result of Wang and Wu [10], $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has average degree at least $2k_{i}$. Now use Theorem On highly edge-connected subgraphs to obtain disjoint $(k_{1}+1)$- and $(k_{2}+1)$-connected subgraphs with vertex number greater than $2k_{1}$ and $2k_{2}$, respectively. Since Thomassen’s argument also works for edge-connectedness, we may proceed as before.
  • (iv)
    By Theorem On highly edge-connected subgraphs and Theorem On highly edge-connected subgraphs for small $k$, there exists a function $g(k)=\big(1+o(1)\big)k$$g(k)\leq 2k$ such that for every $k$, every graph of minimum degree at least $g(k)$ contains a $(k+1)$-edge-connected subgraph. We choose $f(r)=\max\big\{g(r-k)+g(k)+1\,\big|\,1\leq k\leq\frac{r}{2}\big\}$. We have that $f(r)=\big(1+o(1)\big)r$ for $r\to\infty$:
    If $k\geq\sqrt{r}$, then $g(r-k)+g(k)=\big(1+o(1)\big)(r-k)+\big(1+o(1)\big)k=\big(1+o(1)\big)r$.Otherwise, $g(r-k)+g(k)<\big(1+o(1)\big)(r-k)+2k\leq\big(1+o(1)\big)r+o(r)=\big(1+o(1)\big)r$.
    Now let $G$ be a $(k_{1}+k_{2}+1)$-edge-connected graph of minimum degree at least $f(k_{1}+k_{2})\geq g(k_{1})+g(k_{2})+1$. Again by Stiebitz [8], $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has minimum degree at least $g(k_{i})$. We obtain disjoint $(k_{1}+1)$- and $(k_{2}+1)$-edge-connected subgraphs. Then proceed as before. $\Box$
The result of Wang and Wu [10] may also be applied to obtain average degree versions of (i) and (ii).